OpenStudy (anonymous):

Help,Please! Albert and Paul are on the school baseball team. Albert has a batting average of .4, and Paul has a batting average of .3. That means that Albert gets a hit 40% of his at bats and Paul gets a hit 30% of his times at bat. What is the probability that—

OpenStudy (anonymous):

1. both Albert and Paul are able to get hits their first time at bat? (.4/100)*(.3/100) 2) . neither Albert nor Paul is able to get a hit their first time at bat? 100-(40+300=30 =30/100 ? 3) at least one of the two friends is able to get a hit their first time at bat?

OpenStudy (a_clan):

1. 0.4*0.3=0.12 2. (1-0.4)*(1-0.3) 3. 1- {(1-0.4)*(1-0.3)}